# Phase Frustration Induced Intrinsic Bose Glass in the Kitaev-Bose-Hubbard Model

### v1

@@ VI Discussion
The emergence of the Bubble Phase in our frustrated Hamiltonian is not an isolated phenomenon but rather a specific manifestation of a generic mechanism driving frustration-induced localization. While the coexistence of superfluid clusters and insulating barriers is conventionally attributed to quenched disorder [2, 4, 5], our work demonstrates that it can be generically understood through a unified framework of phase frustration. This framework bridges the gap between disorder-driven phenomena and similar fragmented states arising from deterministic inhomogeneities, such as tunable weak links [26], off-diagonal confinement [27], twisted bilayer geometries [12], and even non-Hermitian models characterized by non-reciprocal phase factors [28]. To demystify the physical origin of this frustration, let us consider a generic bosonic system where the hopping amplitude $t_{ij}$ is spatially modulated. Since the hopping connects two sites, its magnitude is naturally defined on the bond center $\bm{R}_{ij}=(\bm{r}_{i}+\bm{r}_{j})/2$. Any such spatial texture $t(\bm{R}_{ij})$ can be decomposed via Fourier analysis into a spectrum of spatial frequencies:
$t_{ij}=t(\bm{R}_{ij})=\sum_{\bm{q}}\mathcal{T}_{\bm{q}}\,e^{i\bm{q}\cdot\bm{R}_{ij}}.$ (6)
This decomposition offers a direct mapping to the standard Peierls substitution, $t_{ij}\sim e^{i\bm{A}(\bm{R}_{ij})\cdot(\bm{r}_{i}-\bm{r}_{j})}$, describing bosons in a gauge field. To see this explicitly, we define the bond vector $\bm{\delta}=\bm{r}_{i}-\bm{r}_{j}$ (assuming unit lattice spacing, $|\bm{\delta}|=1$). For a fixed hopping direction $\bm{\delta}$, we can rewrite the spatial modulation phase using the identity $\bm{q}\cdot\bm{R}_{ij}=[(\bm{q}\cdot\bm{R}_{ij})\bm{\delta}]\cdot\bm{\delta}$. This allows us to identify an effective gauge potential:
$\bm{A}_{\text{eff}}(\bm{R}_{ij})\equiv(\bm{q}\cdot\bm{R}_{ij})\,\bm{\delta}.$ (7)
@@
Substituting this back, we recover the Peierls form $\bm{q}\cdot\bm{R}_{ij}=\bm{A}_{\text{eff}}(\bm{R}_{ij})\cdot(\bm{r}_{i}-\bm{r}_{j})$. Since $\bm{A}_{\text{eff}}(\bm{R})$ is a linear function of the coordinates $\bm{R}$, it corresponds physically to a uniform magnetic field (where the vector potential grows linearly). Consequently, Eq. (6) implies that an arbitrary hopping texture involving multiple $\bm{q}$ modes is equivalent to subjecting the boson to a superposition of distinct, competing magnetic fields. The boson is thus simultaneously driven by conflicting Gauge-imposed constraints, leading to severe phase frustration. To resolve this tension, the system spontaneously creates insulating domain walls to screen the conflicting phase gradients, resulting in the localized superfluid clusters (bubbles) we observe.
@@
This perspective offers a fresh insight into the nature of the Bose Glass. Traditionally, the Bose Glass is viewed as a consequence of Anderson localization induced by diagonal disorder (random chemical potentials) [2]. However, in the context of off-diagonal disorder (random hopping amplitudes) [5], the physics is governed by the complex spatial texture of $t_{ij}$. Applying our Fourier analysis framework [Eq. (6)], a stochastic distribution of hopping strengths corresponds to a broad, continuous spectrum of $\mathbf{q}$ modes. Physically, this represents the limit of “extreme phase frustration,” where the boson is subjected to a turbulent superposition of random effective magnetic fields. Unable to satisfy this chaotic set of gauge-imposed constraints, the system’s global phase coherence shatters. To minimize energy, the superfluid fragments into a myriad of disjointed islands separated by insulating boundaries—phenomenologically identical to the state observed in our model. Our “Intrinsic Bose Glass,” generated here by deterministic phase frustration, thus serves as a pristine, disorder-free archetype of this universal mechanism, revealing that severe phase frustration alone is sufficient to drive the superfluid-to-glass transition.
Acknowledgements.
This work utilizes the Energy Penalty Method to resolve numerical instabilities in the BdG spectrum.
@@ Acknowledgements.
In this section, we provide the explicit expressions for the matrix elements of the Bogoliubov-de Gennes (BdG) dynamical matrix $\mathcal{M}_{\text{BdG}}$ derived within the slave-particle formalism.
As introduced in the main text, the fluctuation operators are defined via the slave particles $\hat{c}_{n,i}$, where $\hat{b}_{i}=\sum_{n}\sqrt{n}\hat{c}_{n,i}^{\dagger}\hat{c}_{n-1,i}$. The BdG Hamiltonian in the Nambu basis $\delta\hat{\Phi}=(\delta\hat{\mathbf{c}},\delta\hat{\mathbf{c}}^{\dagger})^{T}$ takes the block form:
@@
Here, the indices of the submatrices run over $(i,m)$ and $(j,n)$, representing the $m$-th Fock state at site $i$ and the $n$-th Fock state at site $j$, respectively. The hopping amplitude is denoted by $J$, the pairing amplitude by $\Delta$, and the Peierls phase by $\theta_{ij}$. We assume the pairing phase follows the gauge field ($\phi_{ij}=\theta_{ij}$).
The mean-field parameters are defined based on the local ground state coefficients $c_{n,i}^{(0)}$:
@@
In the slave-particle formalism, the Hilbert space is enlarged, introducing a set of spurious zero-energy modes arising from the local $U(1)$ gauge symmetry and normalization constraints. These modes satisfy $\mathcal{M}_{\text{BdG}}|\Psi_{\text{spur}}\rangle=0$. To decouple them from the physical spectrum, we construct an effective Hamiltonian $\mathcal{H}_{\text{eff}}=\mathcal{M}_{\text{BdG}}+\Lambda\mathcal{P}_{\text{spur}}$, where $\mathcal{P}_{\text{spur}}$ is the projection operator onto the spurious subspace and $\Lambda$ is a large penalty parameter.
Here, we provide a rigorous justification based on standard perturbation theory (specifically, the projection technique for effective Hamiltonians). We demonstrate that, within a perturbative expansion in powers of $1/\Lambda$, this method effectively isolates physical modes with a vanishingly small energy shift, even in the presence of non-Hermitian couplings.

### v2

@@ Abstract
Discussion—The emergence of the Bubble Phase in our frustrated Hamiltonian is not an isolated phenomenon but rather a specific manifestation of a generic mechanism driving frustration-induced localization. While the coexistence of superfluid clusters and insulating barriers is conventionally attributed to quenched disorder [2, 4, 5], our work demonstrates that it can be generically understood through a unified framework of phase frustration. This framework bridges the gap between disorder-driven phenomena and similar fragmented states arising from deterministic inhomogeneities [28, 29, 12, 30].
@@
To demystify the physical origin of this frustration, consider a generic bosonic system where the hopping amplitude $t_{ij}$ is spatially modulated. Any such spatial texture $t(\bm{R}_{ij})$ can be decomposed via Fourier analysis into a spectrum of spatial frequencies: $t_{ij}=\sum_{\bm{q}}\mathcal{T}_{\bm{q}}\,e^{i\bm{q}\cdot\bm{R}_{ij}}$. This decomposition offers a direct mapping to the standard Peierls substitution, identifying an effective gauge potential $\bm{A}_{\text{eff}}(\bm{R}_{ij})\equiv(\bm{q}\cdot\bm{R}_{ij})\,\bm{\delta}$. Since $\bm{A}_{\text{eff}}(\bm{R})$ is linear in coordinates, it corresponds physically to a uniform magnetic field. Consequently, an arbitrary hopping texture involving multiple $\bm{q}$ modes is equivalent to subjecting the boson to a superposition of distinct, competing magnetic fields, leading to severe phase frustration. To resolve this tension, the system spontaneously creates insulating domain walls to screen the conflicting phase gradients, resulting in the localized bubbles.
@@
This perspective offers a fresh insight into the nature of the Bose Glass caused by off-diagonal disorder. Traditionally, the Bose Glass is viewed as a consequence of Anderson localization induced by diagonal disorder [2]. However, in the context of off-diagonal disorder [5], a stochastic distribution of hopping strengths corresponds to a broad, continuous spectrum of $\mathbf{q}$ modes. Physically, this represents the limit of “extreme phase frustration.” Unable to satisfy this chaotic set of gauge-imposed constraints, the system’s global phase coherence shatters. Our “Intrinsic Bose Glass,” generated here by deterministic phase frustration, thus serves as a pristine, disorder-free archetype of this universal mechanism.
Data Availability Statement
The code and data used to generate the figures in this study are available from the corresponding author upon reasonable request.
@@
In this section, we provide the explicit expressions for the matrix elements of the Bogoliubov-de Gennes (BdG) dynamical matrix $\mathcal{M}_{\text{BdG}}$ derived within the slave-particle formalism. As introduced in the main text, the fluctuation operators are defined via the slave particles $\hat{c}_{n,i}$, where $\hat{b}_{i}=\sum_{n}\sqrt{n}\hat{c}_{n,i}^{\dagger}\hat{c}_{n-1,i}$. The BdG Hamiltonian in the Nambu basis $\delta\hat{\Phi}=(\delta\hat{\mathbf{c}},\delta\hat{\mathbf{c}}^{\dagger})^{T}$ takes the block form:
@@
Here, the indices of the submatrices run over $(i,m)$ and $(j,n)$, representing the $m$-th Fock state at site $i$ and the $n$-th Fock state at site $j$, respectively. The hopping amplitude is denoted by $J$, the pairing amplitude by $\Delta$, and the Peierls phase by $\theta_{ij}$. We assume the pairing phase follows the gauge field ($\phi_{ij}=\theta_{ij}$). The mean-field parameters are defined based on the local ground state coefficients $c_{n,i}^{(0)}$:
@@
In the slave-particle formalism, the Hilbert space is enlarged, introducing a set of spurious zero-energy modes arising from the local $U(1)$ gauge symmetry and normalization constraints. These modes satisfy $\mathcal{M}_{\text{BdG}}|\Psi_{\text{spur}}\rangle=0$. To decouple them from the physical spectrum, we construct an effective Hamiltonian $\mathcal{H}_{\text{eff}}=\mathcal{M}_{\text{BdG}}+\Lambda\mathcal{P}_{\text{spur}}$, where $\mathcal{P}_{\text{spur}}$ is the projection operator onto the spurious subspace and $\Lambda$ is a large penalty parameter. Here, we provide a rigorous justification based on standard perturbation theory (specifically, the projection technique for effective Hamiltonians). We demonstrate that, within a perturbative expansion in powers of $1/\Lambda$, this method effectively isolates physical modes with a vanishingly small energy shift, even in the presence of non-Hermitian couplings.
@@
Appendix C C. Geometric Stability of the Bubble Phase: Suppression of Percolation
The stability of the Bubble Phase relies on a delicate balance between the bubble density and the bubble size. To stabilize a globally insulating yet compressible state, the system must maintain spontaneous spatial fragmentation. This requires that the superfluid islands (bubbles) remain sufficiently separated by insulating domain walls to prevent global percolation. In this section, we provide a detailed visualization of how increasing $\alpha$ or $\Delta$ drives the system towards an unstable superfluid regime by destroying this separation.
(a) Dependence on Magnetic Flux $\alpha$
(b) Dependence on Pairing Amplitude $\Delta$
Figure S1: Evolution of the real-space superfluid order parameter $|\langle b_{i}\rangle|$ demonstrating the percolation mechanism. (a) Flux dependence (fixed $\Delta=0.005,J/U=0.025,\mu/U=0.8$): As the magnetic flux $\alpha$ increases from $\pi/16$ to $\pi$, the density of induced bubbles increases as $N_{b}\propto\alpha$. At high flux (e.g., $\alpha\geq\pi/8$), the bubbles become densely packed, reducing the mean separation and forcing a breakdown of the insulating barriers. (b) Pairing dependence (fixed $\alpha=\pi/16,J/U=0.025,\mu/U=0.8$): As the pairing amplitude $\Delta$ increases from 0 to $0.025$, the spatial extent (radius) of each bubble expands. Even at a sparse flux density, a large $\Delta$ (e.g., $\Delta>0.025$) causes neighboring bubbles to expand and overlap, initiating superfluid percolation. Therefore, the parameter choice used in the main text ($\alpha=\pi/16,\Delta=0.005$) is critical to maintain sparse, compact bubbles that prevent percolation.
C.1 1. Topological Crowding via Flux $\alpha$
The number of superfluid bubbles is topologically determined by the number of magnetic flux quanta piercing the lattice, scaling as $N_{b}\approx(L\alpha/\pi)^{2}$. Figure S1(a) illustrates this ”crowding effect.”
  • At $\alpha=\pi/16$, the bubbles are sparse, separated by thick insulating domain walls, supporting a robust glassy phase.
  • As $\alpha$ increases , the unit cell shrinks and the bubbles are forced into close proximity. The insulating regions are squeezed into thin filaments, drastically increasing the probability of tunneling between islands and leading to an unstable superfluid state.
C.2 2. Bubble Expansion via Pairing $\Delta$
While $\alpha$ sets the center-to-center distance, the pairing amplitude $\Delta$ (along with hopping $J$) determines the effective radius $R_{b}$ of each bubble. Figure S1(b) demonstrates this ”expansion effect” at a fixed low flux ($\alpha=\pi/16$).
  • At $\Delta=0$, the system is in a Mott insulating state.
  • At $\Delta=0.005$ (our chosen value), clear, distinct bubbles form with finite size, but they remain strictly isolated.
  • Increasing $\Delta$ promotes delocalization, causing the superfluid islands to expand spatially. This expansion progressively thins the insulating domain walls separating the bubbles. Once these barriers are effectively bridged, the localized clusters link up to form a global superfluid.
To realize the Intrinsic Bose Glass phase, it is insufficient to merely introduce frustration. One must simultaneously ensure a sparse density (low $\alpha$) and compact size (low $\Delta$) of the superfluid droplets. Our parameters $\alpha=\pi/16$ and $\Delta=0.005$ are specifically tuned to satisfy this condition, ensuring that transport is blocked by spontaneous spatial fragmentation.